Midland Resources has two production departments (Fabrication and Assembly) and three service departments (Engineering, Administration, and Maintenance). During July, the following costs and service department usage ratios were recorded:

Supplying Department Using Department
Engineering Administration Maintenance Fabrication Assembly
Engineering 0 60 % 0 20 % 20 %
Administration 15 % 0 10 % 50 % 25 %
Maintenance 0 30 % 0 15 % 55 %
Direct cost $ 54,000 $ 304,600 $ 50,000 $ 260,000 $ 90,000

  
Required:

Allocate the service department costs to the two operating departments using the reciprocal method. (Do not round intermediate calculations.)

Allocated to:
Costs Fabrication Assembly
Engineering
Administration
Maintenance
Total
Answer #1

Solution:

Service department cost = Direct cost + Allocated costs

Engineering = $54,000 + 15% of administration

Administration = $304,600 + 60% of engineering + 30% of maintenance

Maintenance = $50,000 + 10% of administration

Administration = $304,600 + 60% ($54,000 + 15% of administration) + 30% ($50,000 + 10% of adminstration

Administration = $304,600 + $32,400 + 0.09 Administration + $15,000 + 0.03 Administration

Administration = $352,000 / 0.88 = $400,000

Engineering = $54,000 + 15% of $400,000 = $114,000

Maintenance = $50,000 + 10% of $400,000 = $90,000

Allocated to:
Costs Fabrication Assembly
Engineering $114,000.00 $22,800.00 $22,800.00
Administration $400,000.00 $200,000.00 $100,000.00
Maintenance $90,000.00 $13,500.00 $49,500.00
Total $236,300.00 $172,300.00